Opened 9 years ago

Closed 3 years ago

#351 closed flight processing (fixed)

EUFAR10/02 (AIRES-CZM), 200/2010, Santander

Reported by: jb362 Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: emca

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EUFAR10_02-2010_200_Santander

Data arrived from ARSF via network transfer on 26/07/2010.

Scientific objectives:

1 Development of a reliable tool to estimate coastal water quality indicators by using remote sensing techniques, both airborne and satellite.

1.1. Development of a set of semi-empirical algorithms based on hyperspectral airborne and in-situ data for estimating the concentrations of the water constituents;
1.2. Migration of the developed algorithms into the satellite workflow.

  1. Implementation of an automatic warning system based on satellite imagery processing;
  1. Inclusion of the information obtained by means of remote sensing in the dataset used for the calibration of regional hydrodynamic models.

Priority: a4m - awaiting confirmation

PI: Castillo-Lopez

Sensors: ¶

  • Eagle (21/12/2010)
  • Hawk (02/02/2011)
  • Leica LIDAR (21/12/2010)
  • RCD Digital Photography (21/12/2010)

Change History (46)

comment:1 Changed 9 years ago by jb362

Awaiting logsheet

comment:2 Changed 9 years ago by jb362

  • Priority changed from immediate to alpha 4 high

comment:3 Changed 9 years ago by ella

  • Description modified (diff)
  • Owner set to ella
  • Status changed from new to accepted

comment:4 Changed 9 years ago by ella

Beginning navigation processing

comment:5 Changed 9 years ago by anch

Starting Applanix (and probably IPAS) processing for this project.

comment:6 Changed 9 years ago by anch

  • Owner changed from ella to anch

comment:7 Changed 9 years ago by iopa

  • Description modified (diff)

Taking camera.

comment:8 Changed 9 years ago by emca

  • Other processors set to emca

Starting LiDAR processing

comment:9 Changed 9 years ago by emca

  • Other processors emca deleted

There is a problem with the IPAS for this project. Moving on to a different project until this can be resolved.

comment:10 Changed 9 years ago by jb362

There is also a problem with the applanix for this project which I do not know how to resolve. Because there is no IPAS .sol file or applanix .sbet file, none of the camera, Lidar, or hyperspectral data can be processed at present.

comment:11 Changed 9 years ago by knpa

  • Description modified (diff)

Data moved from overflow to repository.
Can now be r-synced.

comment:12 Changed 9 years ago by benj

Have downloaded 1s RINEX data from University of Cantabria and copied to repository. Relevant section of email from PI was as follows:

SANTANDER and TORRELAVEGA stations are in relation with the calibration site "Santander1" and the LAREDO station for the area of validation (Santander 2). Santander 1 data taken on 18th July and Santander2  data taken on 19th July.
In the corresponding tabs appear links to download RINEX data.
 
LINKS of interest:
 
http://www.gnss.unican.es/estaciones/cant.asp
 
http://www.gnss.unican.es/estaciones/trvg.asp
 
http://www.gnss.unican.es/estaciones/lare.asp
 
ftp://ftpgnss.digteg.unican.es/rinex
 
Access to real time corrections using the standard protocol NTRIP (Networked Transport of RTCM vía Internet Protocol):
 
URL 	http://ntrip.unican.es:2101/	 
IP:    	193.144.202.127	 
Port: 	21 
User: 	anonimous
 
link: http://www.gnss.unican.es/servicios/services.asp

comment:13 Changed 9 years ago by benj

Note I believe the Santander station mentioned with code CANT is the data I have downloaded from REIN. This is because if you go to the first link given the title of the web page is "Estación de Reinosa"

comment:14 Changed 9 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:15 Changed 9 years ago by iopa

Applanix and IPAS have been completed.

comment:16 Changed 9 years ago by iopa

Camera is ready to be checked.

Test images were images 1-8.

On a different note; eagle/hawk files were named 200a for no apparent reason, the logsheet too does not have a flight letter, so they were renamed with the 'a' removed.

comment:17 Changed 9 years ago by emca

  • Other processors set to emca

Starting LiDAR and Eagle

comment:18 Changed 9 years ago by benj

  • Description modified (diff)

comment:19 Changed 9 years ago by emca

Eagle data is ready for delivery check at
~airborne/workspace/EUFAR10_02-2010_200_Santander/delivery/20101125/

comment:20 Changed 9 years ago by emca

Delivery checked the camera data at
~airborne/workspace/EUFAR10_02-2010_200_Santander/delivery/20101109

Data is good, ready to deliver.

comment:21 Changed 9 years ago by emca

LiDAR data is ready for delivery check at
~airborne/workspace/EUFAR10_02-2010_200_Santander/delivery/20101126

comment:22 Changed 9 years ago by emca

SCT (and GPT) offsets for Eagle data:

Line Number SCT (GPT)
01 +0.03
02 0.00
03 -0.10 (-0.10)
04 0.00
05 0.00
06 -0.02
07 -0.12 (-0.08)
08 +0.03
09 +0.02
10 -0.08 (-0.12)
11 -0.01
12 +0.37

Notes on Eagle processing:

Line 05 had an incorrect GPS stop time in the raw header file so this was changed (original file kept as .hdr-original)
Line 08 had a bend in it so processed using -l 0 10100 in azspec (excluding the last ~400)
Line 12 had no sync information

comment:23 Changed 9 years ago by knpa

Delivery Checking LiDAR

comment:24 Changed 9 years ago by knpa

Delivery check complete.

All looks good and ready to deliver apart from:

  • EUFAR10_02-2010_200.dem should be EUFAR10_02-2010-200.dem
  • Low min Z values (~250m), is this due to water?
  • Readme says DEM resolution is 2m, actually 5m.

comment:25 Changed 9 years ago by benj

  • Summary changed from EUFAR10/01, 200/2010, Santander to EUFAR10/02, 200/2010, Santander

Fixed project ID in ticket (was down as EUFAR10/01)

comment:26 Changed 9 years ago by anch

  • Description modified (diff)

Eagle, Camera, and LiDAR all delivered 21/12/2010

comment:27 Changed 9 years ago by emca

Hawk data is ready for delivery check at
~airborne/workspace/EUFAR10_02-2010_200_Santander/delivery/20110107/

Hawk SCT offsets:

line number SCT
01 -0.01
02 -0.01
03 0.00
04 0.00
05 +0.02
06 0.00
07 0.00
08 -0.01
09 0.01
10 0.00
11 -0.02
12 0.54

Notes: Line 08 contained a bend and was processed with -l 0 10120 in azspec. No sync for line 12. SCT's were difficult to correct as lines are mainly over water and very little overlap (just a partial overlap with a cross line). Used LIDAR and Eagle data to help find SCTs.

comment:28 Changed 8 years ago by anch

Delivery checking Hawk.

comment:29 Changed 8 years ago by anch

Delivery check complete.

All fine.

Perhaps rename DEM to match the AZ command in the readme.

Also there are some overflows over the beach/white topped houses on some lev1 files - perhaps mention this in data quality section.

comment:30 Changed 8 years ago by emca

  • Description modified (diff)

Updated readme as per above.
Hawk data sent to PI 02/02/2011

comment:31 Changed 8 years ago by anch

  • Summary changed from EUFAR10/02, 200/2010, Santander to EUFAR10/02 (AIRES-CZM), 200/2010, Santander

comment:32 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from anch to knpa
  • Status changed from accepted to assigned

Project has been r-synced. Beginning archiving.

comment:33 Changed 8 years ago by knpa

tabi data present.

comment:34 Changed 8 years ago by knpa

applanix/gps contents missing.

Pospac project file missing.

comment:35 Changed 8 years ago by knpa

Tarball created. Made available to NEODC.

comment:36 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion

Backed up by NEODC. Moved to archive disk.

comment:37 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:38 Changed 7 years ago by emca

  • Resolution fixed deleted
  • Status changed from closed to reopened

comment:39 Changed 7 years ago by emca

Reprocessed data using apl software

New (apl) sct values for the following Eagle lines:
line 3: -0.06
line 7: -0.10
line 10: -0.01

comment:40 Changed 7 years ago by emca

Hyperspectral data ready for delivery check

comment:41 Changed 7 years ago by adbe

Hyperspectral reprocessing delivery check:

  • Screenshots are at the end of the ReadMe (remove [htb])
  • DEM name is different from example command
  • Pixel Size should really be around 5
  • Should mention dropped scans in the ReadMe
  • e200043b_mapped_utm30n.bil & e200093b_mapped_utm30n.bil & e200123b_mapped_utm30n.bil & h200013b_mapped_utm30n.bil & h200053b_mapped_utm30n.bil & h200073b_mapped_utm30n.bil & h200093b_mapped_utm30n.bil & h200113b_mapped_utm30n.bil: filesize does not match with hdr.

comment:42 Changed 7 years ago by emca

Made updates to readme.
DEM name matches example command (doesn't match config, as it was manually changed afterwards).
Reprocessed Eagle 4, 9, 12 and Hawk 5 7 9 - problem relates to an apl bug that was fixed ~end may but these line were processed before update.

Adbe - can you check these reprocessed lines again?

comment:43 Changed 7 years ago by adbe

Mapped files load ok.
One named incorrectly: h200073b_mapped_utm30n.bil.bil
Remove the "wrong" directory.

comment:44 Changed 7 years ago by emca

Renamed files and deleted 'wrong' directory. Zipping remaining mapped files.

Reprocessing now complete, data not requested by PI. Ready for archiving.

comment:45 Changed 7 years ago by adbe

Delivered reprocessed Hyperspectral to Elena Castillo-Lopez on 16/08/2012 on disk 126.

comment:46 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from reopened to closed
Note: See TracTickets for help on using tickets.